use undetermined coefficients to find the particular solution to

y''+5y'+4y=7t^2+5t+5
Answer #1

Sol ? y+sy+ 4y = 7++57 +5 -0 yp = At²+ Bt te Jp = QAT qat tB 2A yp put in eqn 0 2A + loat +58 + 4at+ 48t+4C = 7t+stts we